1-Which of the following result from the expectations theory of the yield curve?
I. The observed long-term rate includes a risk premium
II. Long term rates are a function of expected future short term rates
III. An upward slope means that the market is expecting higher future short term rates
IV. The observed yield curve is above the pure expectations yield curve.

a) I only b) I and II only c) II and III only d) II, III and IV only

2-The current three year interest rate is 5.45% and the current two year interest rate is 5.15%. The implied forward rate from year two to year three is ______.

a) 5.30% b) 5.65% c) 6.05% d) 5.84%
Answer #1

Expectation theory is used to predict short term interest rates based on the long term interest rates. It states that investors should get the same return if he invests in 2 years bonds as compared to investing consecutively 2 times in 1 year bond.

1. is wrong as risk premium is not involved in expectation thoery

2. is wrong as short term rates are function of long term rates and not vice versa

3. is correct as upword slope means market is expecting higher short term rates.

4. is wrong as observed yield curve is NOT above the pure expectations yield curve.

2) LEt the implied forward rate is r, then

(1+5.15%)^2*(1+r) = (1+5.45%)^3

i.e. 1.10565225 * (1+r) = 1.172572628625

1+r = 1.060525

r = 6.05%

Hence option C is the correct answer.